OpenStudy (anonymous):

Write the complex number in Trig form r(cos+isin), with the interval [0deg, 360deg) -4+4isqrt3

OpenStudy (anonymous):

\[\begin{cases}r\cos\theta=-4\\ r\sin\theta=4\sqrt3\end{cases}\] \[r^2=r^2(\cos^2\theta+\sin^2\theta)=(-4)^2+(4\sqrt3)^2~~\iff~~r^2=64~~\iff~~r=8\] So then, \[r\cos\theta=-4~~\iff~~8\cos\theta=-4~~\iff~~\cos\theta=-\frac{1}{2}\] and \[r\sin\theta=4\sqrt3~~\iff~~8\sin\theta=4\sqrt3~~\iff~~\sin\theta=\frac{\sqrt3}{2}\] What angle \(\theta\) satisfies these ratios?

OpenStudy (anonymous):

120?

OpenStudy (anonymous):

so i would write 120deg -1/2 sqrt3/2?

OpenStudy (anonymous):

You're right about the angle, but the trig form would be \(r(\cos\theta+i\sin\theta\), or \(8\cos120^\circ+8i\sin120^\circ\).

OpenStudy (anonymous):

and that's just the answer?